OpenStudy (anonymous):

solve: 5(x-6)+9=5x+9

OpenStudy (anonymous):

So we have: \[ 5(x-6)+9=5x+9 \]Distributing: \[ 5x-6(5)+9=5x+9\implies\\ 5x-30+9=5x+9\implies\\ 5x-30=5x\implies\\ -30=0 \]So, since this is not true, the problem has no solutions.

OpenStudy (anonymous):

thanks ! @LolWolf

OpenStudy (anonymous):

Sure thing